# Bi-Hölder invariants in o-minimal structures

### v1

@@ 1. Introduction
More recently, Sampaio [Sampaio2025] established a interesting rigidity phenomenon in the complex category: if a complex analytic germ $(X,0)\subset\mathbb{C}^{n}$ is strongly bi-$\alpha$-Hölder equivalent to the germ $(\mathbb{R}^{k},0)$ of a Euclidean plane, then $(X,0)$ must be smooth at the origin.
Motivated by these developments, we consider the following natural questions for definable germs $(X,0)$ and $(Y,0)$ in an o-minimal structure that are strongly bi-$\alpha$-Hölder equivalent:
@@
    for $r$ sufficiently small. The link is well-defined (up to definable homeomorphism) for all sufficiently small $r>0$. When the choice of radius is irrelevant, we write ${\rm Lk}(X,x_{0})$.
@@
  • ${\rm vol}_{k}$ ($k=0,\dots,n$) denotes the $k$-dimensional Lebesgue (Hausdorff) measure on $\mathbb{R}^{n}$. Every definable set is Lebesgue measurable.
@@ 2.2. O-minimal structures
A set $A\subset\mathbb{R}^{n}$ is called definable if $A$ belongs to $\mathcal{D}_{n}$. A map $f:A\to\mathbb{R}^{m}$ is called definable if its graph is a definable set.
@@
    If $k=1$ and $Y_{i}=(a,b)$, then for every $\varepsilon_{1}<\varepsilon_{2}$ in $(a,b)$, the trivialization $h$ induces strong deformation retracts from $p^{-1}([c,\varepsilon_{2}])$ onto $p^{-1}([c,\varepsilon_{1}])$ for every $c<\varepsilon_{1}$ and from $p^{-1}([\varepsilon_{1},d])$ onto $p^{-1}([\varepsilon_{2},d])$ for every $d>\varepsilon_{2}$. To see this, begin with the strong deformation retract $\Phi:[0,1]\times[c,\varepsilon_{2}]\to[c,\varepsilon_{2}]$ from $c,\varepsilon_{2}]$ onto $c,\varepsilon_{1}]$ defined by
@@ 4. Dimension of tangent cones
$C_{x_{0}}X:=\{v\in\mathbb{R}^{n}:\exists(x_{k})\subset X,(x_{k})\to x_{0},\exists(\lambda_{k})\subset\mathbb{R}_{+},\lambda_{k}x_{k}\to v\}.$
It is known that if $X$ is definable then $C_{x_{0}}X$ is also definable, and moreover $\dim C_{x_{0}}X\leq\dim(X,x_{0})$.
@@
Next, compare $\operatorname{dist}(s,A)$ with the spherical distance to $S_{A}$. On $S^{n-1}$, in a small neighborhood $U$ of $S_{A}$, the functions $s\mapsto\operatorname{dist}(s,A)$ and $s\mapsto\operatorname{dist}(s,S_{A})$ are bi-Lipschitz equivalent. Therefore, there exist constants $\kappa_{1},\kappa_{2}>0$ and $\rho_{0}>0$ such that, for all $0<\rho\leq\rho_{0}$,
$\{\,s\in S^{n-1}:\ \operatorname{dist}(s,S_{A})\leq\kappa_{1}\,\rho^{\,d-1}\,\}\ \subset\ E_{\rho}\ \subset\ \{\,s\in S^{n-1}:\ \operatorname{dist}(s,S_{A})\leq\kappa_{2}\,\rho^{\,d-1}\,\}.$
@@
$c_{1}\,\varepsilon^{\,n-a}\ \leq\ {\rm vol}_{n-1}\big(\{\,s\in S^{n-1}:\ \operatorname{dist}(s,S_{A})\leq\varepsilon\,\}\big)\leq c_{2}\,\varepsilon^{\,n-a}.$
@@
where $C={\rm Lip}(H){\rm Lip}(H^{-1})$ which is independent of $s$. Here, ${\rm Lip}(H)$ and ${\rm Lip}(H^{-1})$ denote the Lipschitz constants of $H$ and $H^{-1}$ respectively. This shows that the Lipschitz constant of $r_{s}^{-1}$ is $\sim\frac{1}{s}$.
@@
Since $r_{s}^{-1}$ is Lipschitz map with constant $\sim 1/s$, this gives
@@
${\rm vol}_{n}\big(h(X)\cap B(0,r)\big)\lesssim r^{\frac{(\frac{1}{\alpha}+1)(\alpha-1)n}{\alpha}}{\rm vol}_{n}\big(X\cap B(0,r)\big),$
where the constant for $\sim$” is independent of $r$.
@@
Observe that $\rho=\rho(r)\to 0$ as $r\to 0$ (because $h^{-1}$ is continuous at 0); by taking $r$ sufficiently small we may assume $\operatorname{diam}(E)$ is arbitrarily small.
Fix such a small $r$ and put $D:=\operatorname{diam}(E)>0$. Let $\varepsilon>0$ be arbitrary. By outer regularity of Lebesgue measure and compactness of $\overline{E}$, there exists a finite family of closed axis-parallel cubes $\{Q_{i}\}_{i=1}^{N}$ of common side length $t$ with
@@
Each cube $Q_{i}$ has side $t$ and therefore diameter $\operatorname{diam}(Q_{i})=t\sqrt{n}$. Denote
@@
the volume of the unit ball in $\mathbb{R}^{n}$. For each $i$ we estimate the image-measure of $Q_{i}$ as follows. By the Hölder bound,
$\operatorname{diam}\big(h(Q_{i})\big)\leq L\big(\operatorname{diam}Q_{i}\big)^{\alpha}=L(\sqrt{n}\,t)^{\alpha}.$
Since $h(Q_{i})$ is contained in a ball of radius $\tfrac{1}{2}\operatorname{diam}(h(Q_{i}))$,
${\rm vol}_{n}\big(h(Q_{i})\big)\leq\omega_{n}\Big(\frac{\operatorname{diam}(h(Q_{i}))}{2}\Big)^{n}\leq\omega_{n}2^{-n}L^{n}n^{\frac{\alpha n}{2}}t^{\alpha n}.$
@@
(4) ${\rm vol}_{n}\big(h(X)\cap B(0,r)\big)\leq C^{\prime}\,\rho^{(\alpha-1)n}{\rm vol}_{n}\big(X\cap B(0,\rho)\big).$
To finish the proof, we have to compare ${\rm vol}_{n}\big(X\cap B(0,\rho)\big)$ and ${\rm vol}_{n}\big(X\cap B(0,r)\big)$. Note from the definition that $(1/L)r^{1/\alpha}\leq\rho\leq Lr^{\alpha}$. Since $\alpha-1<0$,
@@
${\rm vol}_{n}\big(h(X)\cap B(0,r)\big)\lesssim r^{\frac{(\alpha-1)n}{\alpha}}{\rm vol}_{n}\big(X\cap B(0,r)\big)\lesssim r^{\frac{(\frac{1}{\alpha}+1)(\alpha-1)n}{\alpha}}{\rm vol}_{n}\big(X\cap B(0,r)\big).$
Case 2: $r<\rho$. By Lemma 4.10,
@@
${\rm vol}_{n}\big(h(X)\cap B(0,r)\big)\lesssim r^{\frac{(\frac{1}{\alpha}+1)(\alpha-1)n}{\alpha}}{\rm vol}_{n}\big(X\cap B(0,r)\big).$
@@
Since $X$ is definable, by Lemma 4.6, $X$ and $C_{0}X$ is ST-equivalent. That is, there are $d_{1},d_{2}>1$ and $C_{1},C_{2}>0$ such that $X\subset ST_{d_{1}}(C_{0}(X),C_{1})$ and $C_{0}(X)\subset ST_{d_{2}}(X,C_{2})$. Applying Lemma 4.5, we get
@@
(8) ${\rm vol}_{n}\bigg(\varphi(ST_{d\alpha^{2}}(E,C_{5}))\cap B(0,r)\bigg)\lesssim r^{\frac{\big(\frac{1}{\alpha}+1\big)(\alpha-1)n}{\alpha}}{\rm vol}_{n}\bigg(ST_{d\alpha^{2}}(E,C_{5})\cap B(0,r)\bigg).$
@@
$K=K(\alpha):=\frac{\big(\frac{1}{\alpha}+1\big)(\alpha-1)n}{\alpha}+(\alpha^{2}-1)d(n-a).$
@@
$|h(x)-h(x^{\prime})|\leq C\,|x-x^{\prime}|^{\alpha},\qquad|h^{-1}(y)-h^{-1}(y^{\prime})|\leq H\,|y-y^{\prime}|^{\alpha}.$
By the McShane–Whitney extension for Hölder functions (see for example [McShane1934]), each coordinate function $h_{i}:X\to\mathbb{R}$ admits an $\alpha$-Hölder extension $\widetilde{h}_{i}:\mathbb{R}^{n}\to\mathbb{R}$ with the same exponent and a controlled constant; likewise each $k_{j}:Y\to\mathbb{R}$ extends to $\widetilde{k}_{j}:\mathbb{R}^{n}\to\mathbb{R}$. Set $\widetilde{h}=(\widetilde{h}_{1},\dots,\widetilde{h}_{n}):\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $\widetilde{k}=(\widetilde{k}_{1},\dots,\widetilde{k}_{n}):\mathbb{R}^{n}\to\mathbb{R}^{n}$. Then $\widetilde{h}|_{X}=h$ and $\widetilde{k}|_{Y}=h^{-1}$, and there are constants $C^{\prime},H^{\prime}>0$ such that
$|\widetilde{h}(u)-\widetilde{h}(u^{\prime})|\leq C^{\prime}|u-u^{\prime}|^{\alpha},\qquad|\widetilde{k}(v)-\widetilde{k}(v^{\prime})|\leq H^{\prime}|v-v^{\prime}|^{\alpha}\quad(u,u^{\prime},v,v^{\prime}\in\mathbb{R}^{n}).$
@@
$\bigl|\,\Phi_{2}(u,v)-\Phi_{2}(u^{\prime},v^{\prime})\,\bigr|$ $=\bigl|\,v+\widetilde{h}(u)-v^{\prime}-\widetilde{h}(u^{\prime})\,\bigr|\ \leq\ |v-v^{\prime}|+C^{\prime}|u-u^{\prime}|^{\alpha},$
$\bigl|\,\Phi_{1}(u,v)-\Phi_{1}(u^{\prime},v^{\prime})\,\bigr|$ $=\bigl|\,u-u^{\prime}-\widetilde{k}(v+\widetilde{h}(u))+\widetilde{k}(v^{\prime}+\widetilde{h}(u^{\prime}))\,\bigr|$
$\leq|u-u^{\prime}|+H^{\prime}\,\bigl|\,(v-v^{\prime})+(\widetilde{h}(u)-\widetilde{h}(u^{\prime}))\,\bigr|^{\alpha}$
$\leq|u-u^{\prime}|+H^{\prime}\bigl(|v-v^{\prime}|+C^{\prime}|u-u^{\prime}|^{\alpha}\bigr)^{\alpha}$
$\lesssim|u-u^{\prime}|+|v-v^{\prime}|^{\alpha}+|u-u^{\prime}|^{\alpha^{2}}.$
$\lesssim\|(u,v)-(u^{\prime},v^{\prime})\|^{\alpha^{2}}\text{ since }\alpha\in(0,1].$
The same argument applied to $\Psi$ (symmetry of the formulas) shows that $\Psi$ is $\alpha^{2}$-Hölder. Therefore $\Phi$ is bi-$\alpha$-Hölder.
@@ 5. Homotopy types of the links of the tangent cones
Let $S_{A}:=A\cap S^{n-1}$ and define the definable function
@@
Consider the definable set
$\Gamma:=\{(t,u)\in[0,\infty)\times S^{n-1}:\ g(u)=t\},$
and the projection $\rho:\Gamma\to(0,\infty)$, $\rho(t,u)=t$. By Hardt’s triviality theorem, there exists $\delta>0$ such that $\rho$ is a definable trivial over $(0,\delta)$. Consequently, for $0<\varepsilon_{1}<\varepsilon_{2}<\delta$, there is a strong deformation retract from $g^{-1}([0,\varepsilon_{2}]$ onto $g^{-1}([0,\varepsilon_{1}])$ (see Observation 2.9), hence the inclusion $g^{-1}([0,\varepsilon_{1}])\hookrightarrow g^{-1}([0,\varepsilon_{2}])$ induces isomorphisms on all homotopy groups.
Since $S_{A}$ is a compact definable set, it is triangulable; thus there exist arbitrarily small open neighborhoods $U$ admitting a definable deformation retraction onto $S_{A}$. By continuity of $g$ and compactness of $S_{A}$, we may choose two such neighborhoods $U_{1}\subset U_{2}$ and $0<\varepsilon_{1}<\varepsilon_{2}<\delta$ such that
@@
Consider the following commute diagram induced from inclusion maps:
@@
induces isomorphisms on all homotopy groups for all sufficiently small $r>0$ provided $Cr^{d-1}<\delta$. This completes the proof. ∎
@@
Let $(X,0)$ and $(Y,0)$ be closed definable germs in $\mathbb{R}^{n}$. Then, there exists $\alpha_{0}\in(0,1)$ such that if $(X,0)$ and $(Y,0)$ are embedded bi-$\alpha$-Hölder equivalent for some $\alpha\in[\alpha_{0},1]$, then for any $C>0$ there exists $d_{0}>1$ such that ${\rm Lk}(ST_{d}(X,C),0)$ and ${\rm Lk}(ST_{d}(Y,C),0)$ have the same homotopy type for every $1<d<d_{0}$.
@@
Fix a constant $C>0$. Consider the sets
$\Omega_{X}:=\{(d,x)\in(1,\infty)\times\mathbb{R}^{n}:{\rm dist}(x,X)\leq C\|x\|^{d}\}$
and
$\Omega_{Y}:=\{(d,x)\in(1,\infty)\times\mathbb{R}^{n}:{\rm dist}(x,Y)\leq C\|x\|^{d}\}.$
Note that $\Omega_{X}$ and $\Omega_{Y}$ are not definable in the same o-minimal structure as $X$ and $Y$, but they are definable in an exponential o-minimal expansion of this structure (see [MS2002]). Let $\rho(d,x)=d$. By Hardt’s triviality theorem, there exists $d_{0}>1$ such that the restriction $\rho|_{\Omega_{\star}}$ is definably trivial over $(1,d_{0})$ for $\star\in\{X,Y\}$, and moreover these trivializations preserve the origin.
Observation 2.9 implies that for any $1<d<d^{\prime}<d_{0}$, there exist strong deformation retracts
$\text{from }\quad\rho|_{\Omega_{X}}^{-1}([d,d_{0}])=ST_{d}(X,C)\quad\text{ onto }\quad\rho|_{\Omega_{X}}^{-1}([d^{\prime},d_{0}])=ST_{d^{\prime}}(X,C),$
and similarly
$\text{from }\quad\rho|_{\Omega_{Y}}^{-1}([d,d_{0}])=ST_{d}(Y,C)\quad\text{onto}\quad\rho|_{\Omega_{Y}}^{-1}([d^{\prime},d_{0}])=ST_{d^{\prime}}(Y,C),$
and these deformation retracts fix the origin.
Fix an $d\in(1,d_{0})$. It follows from Lemma 4.5 that there are $C_{1},C_{2}>0$ such that
@@
Set $d_{1}=d_{1}(\alpha):=d/\alpha^{2}+(1-\alpha)>d/\alpha^{2}$, $d_{2}:=d_{2}(\alpha)=d\alpha^{2}-(1-\alpha)<d\alpha^{2}$. There is $\alpha_{0}<1$ close to $1$ such that for all $\alpha_{0}\leq\alpha<1$, both $d_{1}$ and $d_{2}$ are greater than $1$. We have
@@
Choosing $d_{3}=d_{1}/\alpha^{2}+(1-\alpha)>d_{1}/\alpha^{2}$, we have

### v2

@@ 1. Introduction
More recently, Sampaio [Sampaio2025] established a interesting rigidity phenomenon in the complex category: if a complex analytic germ $(X,0)\subset\mathbb{C}^{n}$ is strongly bi-$\alpha$-Hölder equivalent to the germ $(\mathbb{R}^{k},0)$ of a Euclidean plane, then $(X,0)$ must be smooth.
Motivated by these developments, we consider the following natural questions, ordered by decreasing difficulty, for definable germs $(X,0)$ and $(Y,0)$ in an o-minimal structure that are strongly bi-$\alpha$-Hölder equivalent:
@@
    for $r$ sufficiently small. The link is well-defined up to definable homeomorphism. When the choice of radius is irrelevant, we write ${\rm Lk}(X,x_{0})$.
@@
  • ${\rm vol}_{k}$ ($k=0,\dots,n$) denotes the $k$-dimensional Lebesgue (Hausdorff) measure on $\mathbb{R}^{n}$. Note that every definable set is Lebesgue measurable.
@@ 2.2. O-minimal structures
A set in $\mathcal{D}_{n}$ for some $n$ is called definable set. A map $f:A\to\mathbb{R}^{m}$ is called definable if its graph is a definable set.
@@
    If $k=1$ and $Y_{i}=(a,b)$, then for every $\varepsilon_{1}<\varepsilon_{2}$ in $(a,b)$, the trivialization $h$ induces strong deformation retracts from $p^{-1}([c,\varepsilon_{2}])$ onto $p^{-1}([c,\varepsilon_{1}])$ for every $c\leq\varepsilon_{1}$ and from $p^{-1}([\varepsilon_{1},d])$ onto $p^{-1}([\varepsilon_{2},d])$ for every $d\geq\varepsilon_{2}$. To see this, begin with the strong deformation retract $\Phi:[0,1]\times[c,\varepsilon_{2}]\to[c,\varepsilon_{2}]$ from $c,\varepsilon_{2}]$ onto $c,\varepsilon_{1}]$ defined by
@@ 4. Dimension of tangent cones
$C_{x_{0}}(X):=\{v\in\mathbb{R}^{n}:\exists(x_{k})\subset X,(x_{k})\to x_{0},\exists(\lambda_{k})\subset\mathbb{R}_{+},\lambda_{k}x_{k}\to v\}.$
It is known that if $X$ is definable then $C_{x_{0}}X$ is also definable, and moreover $\dim C_{x_{0}}(X)\leq\dim(X,x_{0})$.
Definition 4.3 (Cone with vertex at the origin).
A subset $A\subset\mathbb{R}^{n}$ is called a cone with vertex at the origin (or simply a cone at 0) if for every $x\in A$ and every $t\geq 0$, we have $tx\in A$.
@@
Thus,
$E_{\rho}=\{s\in S^{n-1}:{\rm dist}(s,A)\leq C\rho^{d-1}\}.$
Let us denote $d_{S^{n-1}}(s,S_{A})$ the spherical distance to $S_{A}$. On $S^{n-1}$, in a small neighborhood $U$ of $S_{A}$, the functions ${\rm dist}(s,A)\sim d_{S^{n-1}}(s,S_{A})$. Therefore, there exist constants $\kappa_{1},\kappa_{2}>0$ and $\rho_{0}>0$ such that, for all $0<\rho\leq\rho_{0}$,
$\{\,s\in S^{n-1}:\ d_{S^{n-1}}(s,S_{A})\leq\kappa_{1}\,\rho^{\,d-1}\,\}\ \subset\ E_{\rho}\ \subset\ \{\,s\in S^{n-1}:\ d_{S^{n-1}}(s,S_{A})\leq\kappa_{2}\,\rho^{\,d-1}\,\}.$
@@
$c_{1}\,\varepsilon^{\,n-a}\ \leq\ {\rm vol}_{n-1}\big(\{\,s\in S^{n-1}:\ d_{S^{n-1}}(s,S_{A})\leq\varepsilon\,\}\big)\leq c_{2}\,\varepsilon^{\,n-a}.$
@@
where $C={\rm Lip}(H){\rm Lip}(H^{-1})$ which is independent of $s$. Here, ${\rm Lip}(H)$ and ${\rm Lip}(H^{-1})$ denote the Lipschitz constants of $H$ and $H^{-1}$ respectively. This shows that the Lipschitz constant of $r_{s}^{-1}$ is $\lesssim\frac{1}{s}$.
@@
Since $r_{s}^{-1}$ is Lipschitz map with constant $\lesssim 1/s$, this gives
@@
${\rm vol}_{n}\big(h(X)\cap B(0,r)\big)\lesssim r^{(\frac{1}{\alpha}+1)(\alpha-1)n}{\rm vol}_{n}\big(X\cap B(0,r)\big),$
where the constant for $\lesssim$” is independent of $r$.
@@
Fix a sufficiently small $r$ and put $D:={\rm diam}(E)>0$ where ${\rm diam}(E)$ denotes the diameter of $E$. Let $\varepsilon>0$ be arbitrary. By outer regularity of Lebesgue measure and compactness of $\overline{E}$, there exists a finite family of closed axis-parallel cubes $\{Q_{i}\}_{i=1}^{N}$ of common side length $t$ with
@@
Each cube $Q_{i}$ has side $t$ and therefore ${\rm diam}(Q_{i})=t\sqrt{n}$. Denote
@@
the volume of the unit ball in $\mathbb{R}^{n}$. By the Hölder bound,
${\rm diam}\big(h(Q_{i})\big)\leq L\big({\rm diam}Q_{i}\big)^{\alpha}=L(\sqrt{n}\,t)^{\alpha}.$
By Jung’s theorem, $h(Q_{i})$ is contained in a ball of radius $\sqrt{1/2}{\rm diam}(h(Q_{i}))$,
${\rm vol}_{n}\big(h(Q_{i})\big)\leq\omega_{n}\Big(\frac{{\rm diam}(h(Q_{i}))}{2}\Big)^{n}\leq\omega_{n}2^{-n/2}L^{n}n^{\frac{\alpha n}{2}}t^{\alpha n}.$
@@
(4) ${\rm vol}_{n}\big(h(X)\cap B(0,r)\big)\leq{\rm vol}_{n}(h(E))\leq C^{\prime}\,\rho^{(\alpha-1)n}{\rm vol}_{n}\big(X\cap B(0,\rho)\big).$
To finish the proof, we have to compare ${\rm vol}_{n}\big(X\cap B(0,\rho)\big)$ and ${\rm vol}_{n}\big(X\cap B(0,r)\big)$. Note from the definition that $(1/L)r^{1/\alpha}\leq\rho\leq Lr^{\alpha}$. Since $\alpha-1<0$, the left-hand-side inequality implies that
@@
and the right-hand-side one gives
(6) $\frac{\rho}{r}\leq Lr^{\alpha-1}$
@@
${\rm vol}_{n}\big(h(X)\cap B(0,r)\big)\lesssim r^{\frac{(\alpha-1)n}{\alpha}}{\rm vol}_{n}\big(X\cap B(0,r)\big)\lesssim r^{(\frac{1}{\alpha}+1)(\alpha-1)n}{\rm vol}_{n}\big(X\cap B(0,r)\big).$
Case 2: $r<\rho$. By Lemma 4.11 and (6),
@@
${\rm vol}_{n}\big(h(X)\cap B(0,r)\big)\lesssim r^{(\frac{1}{\alpha}+1)(\alpha-1)n}{\rm vol}_{n}\big(X\cap B(0,r)\big).$
@@
Since $X$ is definable, by Lemma 4.7, $X$ and $C_{0}(X)$ is ST-equivalent. That is, there are $d_{1},d_{2}>1$ and $C_{1},C_{2}>0$ such that $X\subset ST_{d_{1}}(C_{0}(X),C_{1})$ and $C_{0}(X)\subset ST_{d_{2}}(X,C_{2})$. Applying Lemma 4.6, we get
@@
(9) ${\rm vol}_{n}\bigg(\varphi(ST_{d\alpha^{2}}(E,C_{5}))\cap B(0,r)\bigg)\lesssim r^{(\frac{1}{\alpha}+1)(\alpha-1)n}{\rm vol}_{n}\bigg(ST_{d\alpha^{2}}(E,C_{5})\cap B(0,r)\bigg).$
@@
$K=K(\alpha):=(\frac{1}{\alpha}+1)(\alpha-1)n+(\alpha^{2}-1)d(n-a).$
@@
$\|h(x)-h(x^{\prime})\|\leq C\,\|x-x^{\prime}\|^{\alpha},\qquad\|h^{-1}(y)-h^{-1}(y^{\prime})\|\leq H\,\|y-y^{\prime}\|^{\alpha}.$
By the McShane–Whitney extension for Hölder functions (see for example [McShane1934]), each coordinate function $h_{i}:X\to\mathbb{R}$ admits an $\alpha$-Hölder extension $\widetilde{h}_{i}:\mathbb{R}^{n}\to\mathbb{R}$; and similarly each $k_{j}:Y\to\mathbb{R}$ extends to $\widetilde{k}_{j}:\mathbb{R}^{n}\to\mathbb{R}$. Set $\widetilde{h}=(\widetilde{h}_{1},\dots,\widetilde{h}_{n}):\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $\widetilde{k}=(\widetilde{k}_{1},\dots,\widetilde{k}_{n}):\mathbb{R}^{n}\to\mathbb{R}^{n}$. Then $\widetilde{h}|_{X}=h$ and $\widetilde{k}|_{Y}=h^{-1}$, and there are constants $C^{\prime},H^{\prime}>0$ such that
$\|\widetilde{h}(u)-\widetilde{h}(u^{\prime})\|\leq C^{\prime}\|u-u^{\prime}\|^{\alpha},\qquad\|\widetilde{k}(v)-\widetilde{k}(v^{\prime})\|\leq H^{\prime}\|v-v^{\prime}\|^{\alpha}\quad(u,u^{\prime},v,v^{\prime}\in\mathbb{R}^{n}).$
@@
$\bigl\|\,\Phi_{2}(u,v)-\Phi_{2}(u^{\prime},v^{\prime})\,\bigr\|$ $=\bigl\|\,v+\widetilde{h}(u)-v^{\prime}-\widetilde{h}(u^{\prime})\,\bigr\|$
$\leq\ \|v-v^{\prime}\|+C^{\prime}\|u-u^{\prime}\|^{\alpha},\lesssim\|(u,v)-(u^{\prime},v^{\prime})\|^{\alpha^{2}}\text{ since }\alpha\in(0,1];$
and
$\bigl\|\,\Phi_{1}(u,v)-\Phi_{1}(u^{\prime},v^{\prime})\,\bigr\|$ $=\bigl\|\,u-u^{\prime}-\widetilde{k}(v+\widetilde{h}(u))+\widetilde{k}(v^{\prime}+\widetilde{h}(u^{\prime}))\,\bigr\|$
$\leq\|u-u^{\prime}\|+H^{\prime}\,\bigl\|\,(v-v^{\prime})+(\widetilde{h}(u)-\widetilde{h}(u^{\prime}))\,\bigr\|^{\alpha}$
$\leq\|u-u^{\prime}\|+H^{\prime}\bigl(\|v-v^{\prime}\|+C^{\prime}\|u-u^{\prime}\|^{\alpha}\bigr)^{\alpha}$
$\lesssim\|u-u^{\prime}\|+\|v-v^{\prime}\|^{\alpha}+\|u-u^{\prime}\|^{\alpha^{2}}.$
$\lesssim\|(u,v)-(u^{\prime},v^{\prime})\|^{\alpha^{2}}.$
This implies that $\Phi$ is $\alpha^{2}$-Hölder. The same argument applied to $\Psi$ (symmetry of the formulas) shows that $\Psi$ is $\alpha^{2}$-Hölder. Therefore, $\Phi$ is bi-$\alpha^{2}$-Hölder.
@@ 5. Homotopy types of the links of the tangent cones
Let $S_{A}:=A\cap S^{n-1}$ and define the definable continuous function
@@
By Hardt’s triviality theorem, there exists $\delta>0$ such that $g$ is a definably trivial over $(0,\delta)$. Consequently, for $0<\varepsilon_{1}<\varepsilon_{2}<\delta$, there is a strong deformation retract from $g^{-1}([0,\varepsilon_{2}]$ onto $g^{-1}([0,\varepsilon_{1}])$ (see Observation 2.9), hence the inclusion $g^{-1}([0,\varepsilon_{1}])\hookrightarrow g^{-1}([0,\varepsilon_{2}])$ induces isomorphisms on all homotopy groups.
Since $S_{A}$ is a compact definable set, it is triangulable; thus there exist arbitrarily small open neighborhood $U$ admitting a definable deformation retraction onto $S_{A}$. By continuity of $g$ and compactness of $S_{A}$, we may choose two such neighborhoods $U_{1}\subset U_{2}$ and $0<\varepsilon_{1}<\varepsilon_{2}<\delta$ such that
@@
Consider the following commutative diagram induced from inclusion maps:
@@
induces isomorphisms on all homotopy groups for sufficiently small $r>0$ provided $Cr^{d-1}<\delta$. This completes the proof. ∎
@@
Let $(X,0)$ and $(Y,0)$ be closed definable germs in $\mathbb{R}^{n}$. For every $C>0$ there exists $d_{0}>1$ and $\alpha_{0}\in(0,1)$ such that if $(X,0)$ and $(Y,0)$ are embedded bi-$\alpha$-Hölder equivalent for some $\alpha\in[\alpha_{0},1]$, then for every $d\in(1,d_{0})$ the links ${\rm Lk}(ST_{d}(X,C),0)$ and ${\rm Lk}(ST_{d}(Y,C),0)$ have the same homotopy type.
@@
Fix $C>0$. For a definable germ $Z\subset\mathbb{R}^{n}$ (with $Z=X$ or $Z=Y$), set
$\Omega_{Z}:=\{x\in B(0,1):{\rm dist}(x,Z)\leq C\|x\|^{d}\text{ for some }d\in(1,2]\},$
and define
$\rho_{Z}(x):=\sup\{\,d\in(1,2]:{\rm dist}(x,Z)\leq C\|x\|^{d}\,\},\qquad x\in\Omega_{Z}.$
For $x\in\Omega_{Z}\setminus Z$, the function
$q_{Z}(x):=\frac{\log({\rm dist}(x,Z)/C)}{\log\|x\|}$
is continuous. Since
$\rho_{Z}(x)=\min\{2,q_{Z}(x)\},$
$\rho_{Z}$ is continuous on $\Omega_{Z}\setminus Z$.
Although $\Omega_{Z}$ need not be definable in the original structure, it is definable in an exponential o-minimal expansion (see [MS2002]). In particular, $\rho_{Z}$ is definable on the same structure as $\Omega_{Z}$. Applying Hardt’s triviality theorem to $\rho_{Z}|_{\Omega_{Z}\setminus Z}$, we obtain $\delta_{Z}>1$ such that $\rho_{Z}$ is definably trivial over $(1,\delta_{Z})$. By Observation 2.9, for any $1<d<d^{\prime}<\delta_{Z}$, there exist strong deformation retract from $\rho_{Z}^{-1}([d,\delta_{Z}])=ST_{d}(Z,C)\setminus X$ onto $\rho_{Z}^{-1}([d^{\prime},\delta_{Z}])=ST_{d^{\prime}}(Z,C)\setminus Z$. Since $Z\subset ST_{d}^{\prime}(Z,C)$, we may extend this deformation by defining it to be the identity on $Z$. This yields a strong deformation retract from $ST_{d}(Z,C)$ onto $ST_{d^{\prime}}(Z,C)$. This deformation retract fixes $Z$, hence fixes 0 as well. Consequently, it induces a strong deformation retract from $ST_{d}(Z,C)\setminus\{0\}$ onto $ST_{d^{\prime}}(Z,C)\setminus\{0\}$.
Let $\varphi:(\mathbb{R}^{n},0)\to(\mathbb{R}^{n},0)$ be a bi-$\alpha$-Hölder homeomorphism such that $\varphi(X)=Y$. Set $d_{0}:=\min\{\delta_{X},\delta_{Y}\}$ and fix an $d\in(1,d_{0})$. It follows from Lemma 4.6 that there are $C_{1},C_{2}>0$ such that
@@
Set $d_{1}=d_{1}(\alpha):=d/\alpha^{2}+(1-\alpha)>d/\alpha^{2}$, $d_{2}=d_{2}(\alpha):=d\alpha^{2}-(1-\alpha)<d\alpha^{2}$. There is $\alpha_{0}<1$ close to $1$ such that for all $\alpha_{0}\leq\alpha<1$, both $d_{1}$ and $d_{2}$ are greater than $1$. Since $ST_{d_{1}}(Y,C)\subset ST_{d/\alpha^{2}}(Y,C_{1})$ and $ST_{d\alpha^{2}}(Y,C_{2})\subset ST_{d_{2}}(Y,C)$, we have
@@
Choosing $d_{3}=d_{3}(\alpha):=d_{1}/\alpha^{2}+(1-\alpha)>d_{1}/\alpha^{2}$, we have